The tray table on the back of an airplane seat is 7 inches longer than it is wide. Find the length and width of this table if its perimeter is 62 inches.
Answer by ewatrrr(24785)

 
Hi,

P = 2L + 2w = 62in
L + w = 31in
(w+7) + w = 31in
2x = 24in
x = 12in, the width. The Length is 19in
And...Checking
19in + 12in = 31in